OpenStudy (anonymous):

Write an equation for the line that goes through the two points. P(1,2), m =1

OpenStudy (texaschic101):

y = mx + b --- m is your slope and your points (1,2)...x = 1 and y = 2. So basically, the only thing we do not have is b, the y intercept. So we will sub our info into y = mx + b and solve for b. y = mx + b 2 = 1(1) + b 2 = 1 + b 2 - 1 = b 1 = b so your equation is : y = 1x + 1 or just y = x + 1

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

no problem :)